***Please Show Math & Formulas in a legible manner***

Thanks for your help!

  1. Food Lion, reduced the average price on their of Skittles by 1%. In response, sales jumped by 8%.

A.

Calculate the point price elasticity of demand for Skittles

B.

Calculate the optimal price for Skittles if marginal cost is 70¢ per unit.
Answer #1

As per the question

Percentage reduction in the price = 1%

In response to this the sales (quantity demanded) jumped (increased) by 8%

A. Calculate the point price elasticity of demand for Skittles

Point elasticity = % change in Quantity demanded / % change in Price

Point elasticity = 8% / 1% = 8

Elasticity will be 8.

B. Calculate the optimal price for Skittles if marginal cost is 70¢ per unit.

We have to calculate the price and we have given with elasticity and marginal cost. Use the revenue method of calculating the price. Under this method we have to take that at the optimal level of output the marginal cost is equal marginal revenue (MC=MR).

So given marginal cost will be equal to marginal revenue.

Therefore marginal revenue will be 70 per unit

Elasticity calculated above = 8

We have to calculate the price, otherwise called as Average revenue (AR)

According to the revenue method of calculating elasticity

Elasticity = Average Revenue / (Average Revenue – Marginal Revenue)

E = AR/ (AR – MR)

Putting the figures in the above formula

8 = AR / (AR – 70)

8 (AR – 70) = AR

AR = 80

Average revenue or price will be 80

Verification

E = 80 / (80 – 70)

E = 80 / 10

E = 8 (Proved)